For a quantitative trait observed in the F2 generation, the number of

individuals showing one of the extreme phenotypes was 11 out of a

total of 704 F2 progeny. Therefore the number of genes pairs involved

in controlling the trait being studied, under the given set of

experimental conditions is:
Answer #1

(1/4)^n = ratio of extreme phenotypes
       = 11/704
(0.25)^n = 0.0156

solving for n: convert to log
nlog(0.25) = log 0.0156
n(-0.6) = (-1.8)
n = 3
so, 3 gene pairs.